Use ANSI TVM factors whenever possible.

Problem 1

Your company presently uses a crew of its own employees for all major maintenance. On average, 40 major repairs are performed each year at 24 lost production hours per repair. Keeping the crew costs $1,000,000 per year plus any lost production cost. Lost production costs are estimated at $2,500 per hour.

Your company is considering replacing its in-house maintenance work with an outside company (Ajax) working under an annual maintenance contract. This proposal will pay Ajax $2,000,000 per year but your company will still have to pay for all lost production time. Ajax claims that the time per repair will decrease because of the contractors’ substantial maintenance experience. Assuming the number of major repairs and the cost of lost production time remain the same, how much time per repair is the maximum allowable for Ajax to be the preferred alternative?
Answer #1

Answer :

Given

Number of Repairs R=40 per year

Lost production hour per repair L=24 per year

So total production hour loss T=R*L=40*24= 960 hour

Lost cost per hour H=$2500

Crew cost C=$1,000,000

So total repair cost for in house maintenance =C+H*T=1,000,000+2500*960=$3,400,000

If Ajax to be preferred alternative then

Total cost of Ajax should be less than or equal to total repair cost for in house maintenance

Let for Ajex X is lost production per hour

Then total production loss hour = R*X=40X

Annual contract cost for Ajex A =$2,000,000

So total cost for Ajex =A+H*40X=2,000,000+100,000x

Since Total cost for Ajex should be less than or equal to total in house maintenance cost

So 2,000,000+100000X<=$3,400,000

x<=14 hours

So maximum allowable time per repair = 14 hours